If #f(x) =x^(-1/3)#, what is the derivative of the inverse of f(x)?

the answer is #-3x^-4#, but i don't know how to get there.
Answer #1

#d/(dx)(f^-1(x))=-3x^-4#

Explanation:

Here ,

#f(x)=y=x^(-1/3)#

**Let , the function f be one-one and onto.

So , the inverse function of #f(x)# exists.**

#:.y=x^(-1/3)=(1/x)^(1/3)to[because a^-n=1/a^n]#

#:.y^3=((1/x)^(1/3))^3#

#:.y^3=(1/x)^(1/3xx3) to[because(a^m)^n=a^(mn)]#

#:.y^3=1/x#

#:.x=1/y^3#

#:.color(red)(x=y^(-3)#

We know that ,

#f(x)=y=>x=f^-1(y)to[because "definition of inverse function"]#

#:.#:.d/(dx)(f^-1(x))=-3x^-4# x=f^-1(y)=y^-3#

Now changing variable from #yto x,#

#:.color(blue)(f^-1(x)=x^-3#

#:.d/(dx)(f^-1(x))=d/(dx)(x^-3)#

#:.d/(dx)(f^-1(x))=-3x^(-3-1)#

#:.d/(dx)(f^-1(x))=-3x^-4#